please help me solve this.
1.) One number is thrice another number. If their difference is 14, what are the numbers.
2.) andy is seven years older than Mel.ten years ago, he was twice as old as Mel. Find their present ages.
thanks so much!

Answer by rothauserc(4718)

1) let x be one number and 3x the other, then
3x - x = 14
2x = 14
x = 7
the numbers are 7 and 21
2) Andy = Mel + 7
Andy - 10 = 2*(Mel - 10)
substitute for Andy in second equation
Mel + 7 - 10 = 2 * (Mel - 10)
Mel -3 = 2Mel -20
Mel = 17 years old
Andy = 24 years old